OpenStudy (nataliareina):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

note that parallel lines have the same slope... so try to find the slope first. what is the slope of the equation?

OpenStudy (anonymous):

the slope of the equation can be found by putting the equation into slope intercept form which is y=mx+b. m=slope. so, if you change the equation, you get -3y= -4x -12

OpenStudy (anonymous):

your slope is -4

OpenStudy (anonymous):

so slope = ?

OpenStudy (anonymous):

then you can plug in the slope, and the point (-3,4) in 'point-slope form' to find the equation... try :)

OpenStudy (anonymous):

your next equation would be -3(4)= -4(-3) -12 all you have to do is now multiply (:

OpenStudy (texaschic101):

if you put your equation in y = mx + b form, the slope is in the m position 4x - 3y = -12 -- subtract 4x from both sides -3y = -4x - 12 -- divide both sides bt -3 y = 4/3x + 4 slope is 4/3 A parallel line will have the same slope now use y = mx + b slope(m) = 4/3 (-3,4)..x = -3 and y = 4 now we sub 4 = 4/3(-3) + b 4 = - 4 + b 4 + 4 = b 8 = b so your parallel equation is : y = 4/3x + 8

OpenStudy (texaschic101):

unless you need it in standard form.....do you ?